Problem 1. You are given the following data about the economy GDP=6000
YD = 5100
Government deficit = G –T = 200
C = 3800 NX = - 100
Find government expenditure (G), Investment (I), Private Savings (SH ) Note: YD is disposable income
Problem 2. You are given the following data about the economy GDP=6000
IG =800
IN =200
C=4000
G=1100
Government surplus=30
Find net domestic product (NDP), net exports, private savings and depreciation Note: IN are net investments (gros investments net of the depreciation)
Problem 3. Consider a closed economy with expenditure totals given by: C = 1200
I = 400
G = 300
F = 200 (transfers) T=400
a) What is GDP?
b) What is private saving?
c) What is public saving?
d) What is total saving? Show that it equals investment.
Problem 4. Now suppose that the economy from Problem 3 is an open one and that NX=-100. Answer questions a-c. Replace question d) by:
d’) We can define Sr=-NX as the rest of world saving. What is total saving now? Show that it equals investment.
1
Expert's answer
2016-06-02T10:05:03-0400
Problem 1. GDP = 6000, YD = 5100, G –T = 200, C = 3800, NX = -100.
YD = GDP - T, so T = GDP - YD = 6000 - 5100 = 900, then government expenditure G = T + 200 = 900 + 200 = 1100.
GDP = C + I + G + NX, so:.
Investment I = GDP - C - G - NX = 6000 - 3800 - 1100 + 100 = 1200.
Private Savings (SH) equals investment, so SH = I = 1200.

Problem 2. GDP = 6000, IG = 800, IN = 200, C = 4000, G = 1100, G - T = 30.
Net domestic product NDP = GDP - D, where D - depreciation, D = G - IG = 1100 - 800 = 300, so NDP = 6000 - 300 = 5700.
Net exports NX = GDP - C - I - G = 6000 - 4000 - (200 + 300) - (800 + 300) = 400.
Private savings equals investment, so SH = 200 + 300 = 500.

Problem 3. C = 1200, I = 400, G = 300, F = 200 (transfers), T = 400.
a) GDP = C + I + G = 1200 + 400 + 300 = 1900.
b) Private saving is SH = I - F = 400 - 200 = 200.
c) Public saving is SG = T - F = 400 - 200 = 200.
d) Total saving is S = SG + SH = 200 + 200 = 400 = I.

Problem 4. The economy from Problem 3 is an open one, NX = -100.
a) GDP = C + I + G + NX = 1200 + 400 + 300 - 100 = 1800.
b) Private saving is SH = I - F = 400 - 200 = 200.
c) Public saving is SG = T - F = 400 - 200 = 200.
d’) Sr=-NX. Total saving now is S = SG + SH + Sr = 200 + 200 + 100 = 500.